Circular No. 8965

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 P1 (GARRADD) G. J. Garradd reports his discovery of a comet on CCD images ob-

tained in poor seeing with the 0.5-m Uppsala Schmidt telescope (discovery observation tabulated below), the object described as being ∼ 10′′ in size with a faint, broad feature extending ∼ 15′′ to the west. Following posting on the ‘NEOCP’ webpage, C. Jacques and E. Pimentel (Belo Horizonte, Minas Gerais, Brazil, 0.30-m reflector) report that thirty stacked 60-s CCD exposures show a condensed coma of diameter 20′′ on Aug. 14.2 UT.

2008 UT  α_{2000}  δ_{2000}  Mag.
Aug. 13.72362  0^{h}25^{m}59.^{s}88  −46^{o}56′35′′.8  16.7

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2008-P66.

T = 2009 July 2.633 TT  ω = 12.^{o}615
Ω = 0.118 }2000.0
q = 3.64014 AU  i = 61.227

COMET P/2008 O2 (McNAUGHT) Additional astrometry have permitted orbital calculations to show that

this comet (cf. IAUC 8963) is of short period, the elements below taken from MPEC 2008-P54.

T = 2009 Aug. 5.204 TT  ω = 42.^{o}253
e = 0.15655  Ω = 325.529 }2000.0
q = 3.78077 AU  i = 9.967
a = 4.48252 AU n^{o} = 0.103853 P = 9.5 years

COMET P/2008 O3 (BOATTINI) Additional astrometry have permitted orbital calculations to show that

this comet (cf. IAUC 8964) is of short period, the elements below taken from MPEC 2008-P65.

T = 2008 June 4.806 TT  ω = 341.^{o}527
e = 0.72567  Ω = 47.685 }2000.0
q = 2.51228 AU  i = 32.218
a = 9.15770 AU n^{o} = 0.035565 P = 27.7 years
2008 August 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green